Write an equation in terms of x that determines how much water should be added to x liters of an 18% acid solution to reduce the concentration to 15%
Answer by richwmiller(17219)

from view of acid
x*0.18=0.15*(x+y)
or from view of water
1*y+0.82*x=0.85(x+y)
y is acid x is water